In December 2016, Boise Cascade's stock had a beta of 0.97. The Treasury bill rate at the time was 4.34% percent, and the Treasury bond rate was 7.59% percent. The firm had debt outstanding of 1.586 billion and a market value of equity of 2.616 billion; the corporate marginal tax rate was 36.00%. Use 8.00%as the market risk premium. How much of the risk (in percentage terms) in the company can be attributed to business risk and how much to financial leverage risk?

69% and 31%

31% and 69%

72% and 28%

28% and 72%

97% and 3%
Answer #1

Unlevered beta = Beta / (1 +(1- tax rate) (debt / Equity))

= 0.97 / (1 +(1- 36%) (1.586 / 2.616))

=69.9%

Financial risk = 100 - 69% = 31%

Option ,31% and 69%